# Graph Convolutions Enrich the Self-Attention in Transformers!

### v2

@@ 3.2 Theoretical Analysis
In this subsection, we provide a theorem that provides an upper bound on the error introduced by approximating the power of a matrix, specifically using the first-order Taylor expansion. The theorem specifically analyzes the error for the matrix $\bar{\bm{A}}^{K}$ when approximated using a first-order Taylor expansion.
Theorem 3.1 (Error bound for first-order Taylor approximation).
Define the error term as the difference between the exact value and the first-order Taylor approximation of $\bar{\bm{A}}^{K}$ :
${E}_{K}=||\bar{\bm{A}}^{K}-(\bar{\bm{A}}+(K-1)(\bar{\bm{A}}^{2}-\bar{\bm{A}}))||,$ (7)
where |||||$\cdot$||| | ⋅ | | denotes the $L_{\infty}$ norm. Then, the error $E_{K}$ can be bounded as follows:
${E}_{K}\leq 2K.$ (8)
The error bound provides an upper limit for the difference between the true value of $\bar{\bm{A}}^{K}$ and its approximation using the given formula. When the row-sum of $\bar{\bm{A}}$ is 1 due to softmax. The proof of Theorem. 3.1 is in Appendix A.
Computing high-order matrix powers can be computationally intensive, especially for large matrices such as the self-attention in natural language processing where the number of tokens (words) can be tens of thousands. The first-order Taylor approximation provides a simpler computation that can significantly reduce the required computational resources and time. Theorem. 3.1 ensures that even with this simplification, the error can be bounded and is predictable.
3.3 Discussion
This subsection provides a discussion on GFSA in terms of the comparison with other models and the mitigation of the oversmoothing problem.
Comparison to Transformers
In the field of computer vision, there have been recent researches on adjusting the frequency response of ViT. HAT (Bai et al., 2022) improves ViT by directly increasing the high-frequency components of images through adversarial training, i.e., modifying inputs. Wang et al. (2022) suggest that they re-balance the low and high-frequency components of the attention map and feature map, respectively (Wang et al., 2022). In the NLP domain, Shi et al. (2022) redesign BERT with multiple fusions in (Shi et al., 2022), inspired by JKNet (Xu et al., 2018).
Comparison to GCNs
Comparisons to GCNs that can be interpreted as graph filters (Kipf & Welling, 2017; Defferrard et al., 2016; Gasteiger et al., 2019) are inevitable. GFSA without a high-order term is analogous to ChebNet (Defferrard et al., 2016) with $K=1$. In addition, GFSA reduces to the vanilla GCN (Kipf & Welling, 2017) when $K=1$, $w_{0}=0$, $w_{1}=1$. GPR-GNN (Chien et al., 2021), which approximates graph convolutions using the monomial basis, is identical to GFSA if our GFSA only considers up to an order of 1 and learns coefficients. When we use only a high-order term and $w_{K}$ is learned to negative value, GFSA can become similar to the reaction-diffusion layer of GREAD (Choi et al., 2023), $\bar{\mathbf{A}}\mathbf{X}+\beta(\bar{\mathbf{A}}-\bar{\mathbf{A}}^{2})$, depending on the higher order terms.
@@ 3.3 Discussion
The key leading to the behaviors of our proposed filter is the coefficients $\{w_{0},w_{1},w_{K}\}$ — note that in the self-attention of Transformers, $w_{0}=w_{K}=0$ and $w_{1}=1$. Since our method can learn any appropriate values for them for a downstream task, it can reduce to low-pass-only, high-pass-only, or combined filters. In particular, our filter reduces to GPR-GNN and GREAD, two popular method preventing the oversmoothing for graph convolutional networks, in certain learned settings on $\{w_{0},w_{1},w_{K}\}$. We use $\bar{\mathbf{A}}^{2}-\bar{\mathbf{A}}$ to approximate $\bar{\mathbf{A}}^{K}$ (see Equation 4). As mentioned earlier, the low/high-pass filter characteristics are decided by $\{w_{0},w_{1},w_{K}\}$. Therefore, the approximated $\bar{\mathbf{A}}^{K}$ by $(K-1)(\bar{\mathbf{A}}^{2}-\bar{\mathbf{A}})$ and the coefficients $\{w_{0},w_{1},w_{K}\}$ can alleviate the oversmoothing problem of the self-attention by constituting an appropriate graph filter for a downstream task.
@@ 4 Experiments
In this section, we demonstrate the effectiveness of GFSA through a series of experiments. These experiments encompass various tasks: i) language understanding and causal language modeling, ii) image classification, iii) graph classification, and iv) code classification. We replace the self-attention of base Transformers in those fields with our GFSA. Our modification adds only tens to hundreds of parameters, which are negligible in comparison with the original size of base models. We summarize the runtime of base models vs. base models with GFSA in Appendix L.
4.1 Experiments on Language Models
To investigate the efficacy of GFSA, we present experiments on natural language understanding and causal language modeling tasks in the subsections below.
4.1.1 Natural Language Understanding
Experimental Setting.
We integrate GFSA into three pre-trained large language models: BERT, ALBERT, and RoBERTa. We evaluate them on the GLUE benchmark, which includes three categories of natural language understanding tasks: i) single-sentence tasks CoLA and SST-2; ii) similarity and paraphrasing tasks MRPC, QQP, and STS-B; iii) natural language inference tasks MNLI, QNLI, and RTE. For the MNLI task, we experiment on both the matched (MNLI-m) and mismatched (MNLI-mm) versions. Following  (Devlin et al., 2019), we report Matthews correlation for CoLA, F1 scores for QQP and MRPC, Spearman correlations for STS-B, and accuracy scores for the other tasks. For each task, we select the best hyperparameters for GFSA, and the other hyperparameters are fixed. We compare our GFSA with ContraNorm (Guo et al., 2023), one of the related methods that address oversmoothing. We finetune ContraNorm with the recommended hyperparameters in the original paper. We initialize with pretrained language model and finetune with added GFSA for 5 epochs. The detailed experimental settings are in Appendix E.1.
@@ Results.
The results are shown in Table 1. When GFSA was plugged into backbones, average performance scores improved across all models over pure backbones. This indicates that GFSA is effective in both large models like BERT and RoBERTa, as well as relatively smaller models like ALBERT. It is worth mentioning that in the case of RoBERTa finetuned on the CoLA dataset, there is a large margin increase from 60.34% to 64.11%, which is a 3.77% improvement with only 144 additional parameters. When compared to ContraNorm, GFSA shows a greater performance improvement on average. Figure 4 in Appendix D shows that these performance enhancements can be attributed to addressing the oversmoothing issue through the designed graph filter.
Table 1: Results comparison on validation set of GLUE tasks. Avg denotes the average performance on all the tasks.
@@
BERTBASE (Devlin et al., 2019) 110M 56.79 93.81 88.70 88.32 88.16 84.96/84.15 91.63 66.06 82.51
BERTBASE + ContraNorm 110M 59.89 93.92 89.88 88.51 88.36 85.11/84.50 91.84 69.31 83.48
ALBERTBASE (Lan et al., 2019) 11M 57.86 92.32 91.80 85.30 90.37 85.37/84.37 91.76 76.90 84.01
ALBERTBASE + ContraNorm 11M 57.45 93.00 92.83 87.78 90.55 85.06/84.57 92.28 78.70 84.69
ALBERTBASE + GFSA 11M 60.21 93.23 92.47 87.79 90.63 85.29/84.92 92.17 78.70 85.05
RoBERTaBASE (Liu et al., 2019) 125M 60.34 94.84 92.28 88.86 89.99 87.94/87.30 92.57 78.70 85.87
RoBERTaBASE + ContraNorm 125M 63.06 95.41 93.17 88.91 90.34 87.88/87.40 92.82 80.51 86.61
RoBERTaBASE + GFSA 125M 64.11 95.41 93.52 89.09 90.35 87.99/87.54 92.97 80.14 86.79
4.1.2 Causal Language Modeling
Experimental Setting.
We also validate the effectiveness of GFSA on causal language modeling problems. We finetune GPT2 on the following three datasets: Penn Treebank (PTB) (Marcus et al., 1993), WikiText-2 and WikiText-103 (Merity et al., 2016). Following the evaluation method in (Yao et al., 2022), we finetune models for 15 epochs with PTB, 4 epochs with WikiText-103, and 10 epochs with WikiText-2, and report the perplexity for sensitivity metric, i.e., higher perplexity means higher sensitivity. The detailed experimental settings are in Appendix F.1.
@@
Table 33 shows the perplexity on PTB, WitiText-2, and WikiText-103. Across all datasets, GPT2 with GFSA consistently outperforms the vanilla GPT2. Our GFSA improves the average perplexity from 18.806 to 18.764. Note that the performance improvements are made with only 144 additional learnable parameters for 12 layers with 12 heads, which are negligible in comparison with the size of GPT2.
Table 2: Results comparison on GPT-2 finetuned with GFSA
Method #Params PTB WikiText-2 WikiText-103 Avg
GPT2 (Radford et al., 2019) 117M 19.513 20.966 15.939 18.806
GPT2 + GFSA 117M 19.450 20.923 15.919 18.764
@@ 4.2 Experiments on Vision Transformers
Experimental Setting.
We aim to demonstrate the efficacy of our GFSA across a spectrum of ViT backbones with different depth settings and training modes. We choose DeiT (Touvron et al., 2021a), CaiT (Touvron et al., 2021b), and Swin (Liu et al., 2021) as backbones, and models are trained from scratch. When training the 12-layer DeiT, we follow the same training recipe, hyperparameters, and data augmentation from (Touvron et al., 2021a). When training the 24-layer DeiT, we follow the settings of (Gong et al., 2021) and set the dropout rate to 0.2. For Cait, we apply our proposed GFSA only to the patch embedding layer. For detailed experimental settings, see Appendix G.1.
@@ Results.
All experimental evaluations are summarized in Table 3. We compare various state-of-the-art models on the ImageNet-1k benchmark. We select them from three classes: CNN only, CNN + Transformer, and pure Transformer. In the Transformer category, we only test with lighter models having comparable numbers of parameters, such as ViT-S and DeiT-S. The results show that the proposed GFSA successfully enhances Deit, CaiT, and Swin across all depth settings and training methods. GFSA provides additional parameters less than 72 for 12-layer DeiT while improving the top-1 accuracy by 1.63%. Compared to existing techniques, improvements by GFSA already surpasses DeepViT-24B (0.6%) (Zhou et al., 2021a), LayerScale (0.7%) (Touvron et al., 2021b), LateInsertion (0.6%) (Touvron et al., 2021b), and HAT (Bai et al., 2022) (1.38%). To sum up, we observed that both shallow and deep ViTs can achieve the following benefits from GFSA: i) The filter response plot shows GFSA can preserve higher-frequency representation (cf. Figure 2 (a)) and ii) Figure 2 (b) shows that GFSA can mitigate the increase in the cosine similarity of representation as the layer gets deeper. We also show experiments with the same ContraNorm setting in Appendix G.4.
@@
CNN ResNet-152 (He et al., 2016) 224 152 230M 78.1
DenseNet-201 (Huang et al., 2017) 224 201 77M 77.6
CNN + CVT-21 (Wu et al., 2021) 224 21 32M 82.5
Transformer Refiner (Zhou et al., 2021b) 224 16 86M 81.2
Transformer ViT-S/16 (Dosovitskiy et al., 2021) 224 12 49M 78.1
ViT-B/16 (Dosovitskiy et al., 2021) 224 12 86M 79.8
DeiT-S (Touvron et al., 2021a) 224 12 22M 79.8
DeiT-S Distilled (Touvron et al., 2021a) 224 12 22M 81.2
DeiT-S + LayerScale (Touvron et al., 2021b) 224 12 22M 80.5
DeiT-S + LateInsertion (Touvron et al., 2021b) 224 12 22M 80.5
DeiT-S + ClassAttention (Touvron et al., 2021b) 224 12 22M 80.6
DeiT-S + AttnScale (Wang et al., 2022) 224 12 22M 80.7
DeiT-S + FeatScale (Wang et al., 2022) 224 12 22M 80.9
DeiT-S + HAT (Bai et al., 2022) 224 12 22M 80.9
DeiT-S + Diverse (Chen et al., 2022) 224 12 22M 80.6
DeiT-S + ContraNorm (Guo et al., 2023) 224 12 22M 80.4
Swin-S (Liu et al., 2021) 224 12 50M 82.9
T2T-ViT-24 (Yuan et al., 2021) 224 24 64M 82.3
DeepViT-24B (Zhou et al., 2021a) 224 24 36M 80.1
DeiT-S (Touvron et al., 2021a) 224 24 43M 80.5
CaiT-S (Touvron et al., 2021b) 224 24 47M 82.6
DeiT-S + DiversePatch (Gong et al., 2021) 224 24 44M 82.2
DeiT-S + LayerScale (Touvron et al., 2021b) 224 12 22M 82.4
DeiT-S + AttnScale (Wang et al., 2022) 224 24 44M 81.1
DeiT-S + FeatScale (Wang et al., 2022) 224 24 44M 81.3
DeiT-S + ContraNorm (Guo et al., 2023) 224 24 43M 80.7
GFSA DeiT-S + GFSA 224 12 43M 81.1 ($\uparrow$ 1.3)
DeiT-S + GFSA 224 24 43M 81.5 ($\uparrow$ 1.0)
CaiT-S + GFSA 224 24 43M 82.8 ($\uparrow$ 0.2)
Swin-S + GFSA 224 12 50M 83.0 ($\uparrow$ 0.1)

### v5

@@ 4 Properties of GFSA
We provide a theorem that provides an upper bound on the error introduced by approximating the power of a matrix, specifically using the first-order Taylor expansion. The following theorem specifically analyzes the error of matrix $\bar{\bm{A}}^{K}$ when approximated using a first-order Taylor expansion.
Theorem 4.1 (Error bound for approximated high-order term in GFSA).
Define the error term, $E_{K}$, as the difference between the exact value and approximated value of $\bar{\bm{A}}^{K}$, which is given by $E_{K}=||\bar{\bm{A}}^{K}-(\bar{\bm{A}}+(K-1)(\bar{\bm{A}}^{2}-\bar{\bm{A}}))||_{F}$, where $||\cdot||_{F}$ denotes the Frobenius norm. Then, the error bound can be shown that $E_{K}\leq 2\sqrt{n}K.$
The error bound provides an upper limit for the difference between the actual value of $\bar{\bm{A}}^{K}$ and its approximation. The proof of Theorem 4.1 is in Appendix H. It shows theoretical validity for using approximations to the high-order term in the filters of our GFSA. In terms of performance, we report a difference of approximately $\bar{A}^{K}$ between the actual calculated values in Appendix J.
@@
The key leading to the low/high pass filtering behavior of our proposed filter is the coefficients $\{w_{0},w_{1},w_{K}\}$ — note that in the self-attention of Transformers, $w_{0}=w_{K}=0$ and $w_{1}=1$. Since our method can learn any appropriate values for them for a downstream task, it can be reduced to low-pass-only, high-pass-only, or combined filters. According to Theorem 3.1, our graph polynomial filter can be said to be a low-pass filter when $w_{1},w_{K}$ are positive and a high-pass filter when they are negative. Therefore, our method can learn the appropriate coefficients $\{w_{0},w_{1},w_{K}\}$ for downstream tasks, so it can be reduced to a low-pass-only, high-pass-only, or combined filter, alleviating the oversmoothing problem of self-attention.
The meaning of the high-order term in GFSA in the context of Transformers.
Existing self-attention only captures simple pairwise similarities between tokens and is limited in capturing high-order dependencies. For example, given the two sentences, “Books are more expensive than pencils” and “Books are cheaper than computers”, to understand the relationship between “computers” and “pencils”, we need to capture the high-order dependencies connected through the “Book” token. However, it is difficult to capture these high-order dependencies with traditional self-attention [103]. Therefore, from a Transformer perspective, the approximated $\bar{\bm{A}}^{K}$ in GFSA can be interpreted as being able to capture these high-order dependencies.
Comparison to Transformers.
In the field of computer vision, there has been recent research on adjusting the frequency response of ViT. HAT [4] creates adversarial examples by altering clean images with high-frequency perturbations and jointly trains the ViT on clean images and adversarial examples. Through this, they aim to solve the problem of the ViT being unable to capture high-frequency by allowing us to capture the high-frequency components of the images. However, HAT has the disadvantage of requiring more epochs than the existing ViT, as it must perform adversarial training in some initial epochs and train normally in the remaining epochs. Wang et al. [80] use the concept of DSP, which is a special case of GSP, to isolate the lowest frequency component in the Fourier domain and use a filter learned by rescaling the low and high-frequency components. On the other hand, our GFSA extends the concept to graph signal processing and redesigns self-attention as a graph filter. While GFSA seeks to design a better graph filter by interpreting self-attention as a graph filter, Shi et al. [71] are inspired by JKNet [91], and they solve the oversmoothing problem by fusing the hidden vectors of each layer. However, their method has a limitation with memory increasing, and they only applied it to BERT.
Comparison to GCNs.
Comparisons to GCNs that can be interpreted as graph filters [39, 15, 24] are inevitable. GFSA without a high-order term is analogous to ChebNet [15] with $K=1$. In addition, GFSA reduces to the vanilla GCN [39] when $K=1$, $w_{0}=0$, $w_{1}=1$. GPR-GNN [12], which approximates graph convolutions using the monomial basis, is identical to GFSA if it only considers up to first order and additionally uses a $K$-order term and learns the coefficients. When we use only a high-order term and $w_{K}$ is learned to a negative value, GFSA can become similar to the reaction-diffusion layer of GREAD [13], $\bar{\bm{A}}\bm{X}+\beta(\bar{\bm{A}}-\bar{\bm{A}}^{2})$, depending on the higher order terms.
@@ 5 Experiments
In this section, we demonstrate the effectiveness of GFSA through a series of experiments. These experiments encompass various tasks: i) language understanding and causal language modeling, ii) image classification, iii) graph-level task, and iv) code classification. We replace the self-attention of base Transformers in those fields with our GFSA. Our modification adds only tens to hundreds of parameters, which are negligible in comparison with the original size of base models.
5.1 Experiments on Natural Language Understanding
Setting.
We integrate GFSA into 3 pre-trained large language models: BERT, ALBERT, and RoBERTa. We evaluate them on the GLUE benchmark, which includes 3 categories of natural language understanding tasks: i) single-sentence, ii) similarity and paraphrasing, and iii) natural language inference tasks. For each task, we select the best hyperparameters for GFSA, and the other hyperparameters are fixed. The detailed experimental settings are in Appendix K.1.
@@ Results.
The results are shown in Table 1. When GFSA was plugged into backbones, average performance scores improved across all models over pure backbones. This indicates that GFSA is effective in both large models like BERT and RoBERTa, as well as relatively smaller models like ALBERT. It is worth mentioning that in the case of RoBERTa finetuned on the CoLA dataset; there is a significant margin increase from 60.34% to 64.11%, which is a 3.77% improvement with only 144 additional parameters. When compared to ContraNorm, GFSA shows a greater performance improvement on average. Fig. 5 in Appendix C shows that these performance enhancements can be attributed to addressing the oversmoothing issue through the designed graph filter.
Table 1: Results comparison on GLUE benchmark. Avg denotes the average performance.
@@
BERTBASE [16] 110M 56.79 93.81 88.70 88.32 88.16 84.96/84.15 91.63 66.06 82.51
+ ContraNorm 110M 59.89 93.92 89.88 88.51 88.36 85.11/84.50 91.84 69.31 83.48
+ GFSA 110M 59.56 94.15 90.60 88.46 88.33 85.12/85.06 91.95 68.95 83.58
ALBERTBASE [40] 11M 57.86 92.32 91.80 85.30 90.37 85.37/84.37 91.76 76.90 84.01
+ ContraNorm 11M 57.45 93.00 92.83 87.78 90.55 85.06/84.57 92.28 78.70 84.69
+ GFSA 11M 60.21 93.23 92.47 87.79 90.63 85.29/84.92 92.17 78.70 85.05
RoBERTaBASE [44] 125M 60.34 94.84 92.28 88.86 89.99 87.94/87.30 92.57 78.70 85.87
+ ContraNorm 125M 63.06 95.41 93.17 88.91 90.34 87.88/87.40 92.82 80.51 86.61
+ GFSA 125M 64.11 95.41 93.52 89.09 90.35 87.99/87.54 92.97 80.14 86.79
5.2 Experiments on Causal Language Modeling
Setting.
We also validate the effectiveness of GFSA on causal language modeling problems. We finetune GPT2 [61] on the following 3 datasets: Penn Treebank (PTB) [47], WikiText-2, and WikiText-103 [50]. Following the evaluation method in Yao et al. [93], we finetune models for 15 epochs with PTB, 4 epochs with WikiText-103, and 10 epochs with WikiText-2, and report the perplexity for sensitivity metric. The detailed experimental settings are in Appendix L.1.
Table 2: Results comparison on GPT-2 finetuned with GFSA. Avg denotes the average performance.
Method #Params PTB WikiText-2 WikiText-103 Avg
GPT2 117M 19.513 20.966 15.939 18.806
+ GFSA 117M 19.450 20.923 15.919 18.764
@@
Table 2 shows the perplexity on PTB, WitiText-2, and WikiText-103. Across all datasets, GPT2 with GFSA consistently outperforms the vanilla GPT2. Our GFSA improves the average perplexity from 18.806 to 18.764. Note that performance improvements are made with only 144 additional learnable parameters for 12 layers with 12 heads.
@@ 5.3 Experiments on Vision Transformers
Setting.
We aim to demonstrate the efficacy of our GFSA across a spectrum of ViT backbones. We choose DeiT [74], CaiT [75], and Swin [45] as the backbone, and the models are trained from scratch. When training the 12-layer DeiT, we follow the same training recipe, hyperparameters, and data augmentation from Touvron et al. [74]. For detailed experimental settings, see Appendix M.1.
@@ Results.
The experimental evaluations are summarized in Table 3. We compare various models on the ImageNet-1k benchmark. The results show that the proposed GFSA successfully enhances DeiT, CaiT, and Swin across all depth settings and training methods. GFSA provides additional parameters less than 72 for 12-layer DeiT while improving top-1 accuracy by 1.63%. To sum up, we observed that both shallow and deep ViTs can achieve the following benefits from GFSA: i) The filter response shows GFSA can preserve higher-frequency representation (cf. Fig. 2 (a)) and ii) Fig. 2 (b) shows that GFSA mitigates the increase in the cosine similarity of representation as the layer gets deeper. We further compare with state-of-the-art models that use Fourier transforms rather than graph filters in Appendix M.5. We also show results under the same settings as ContraNorm [28] in Appendix M.6.
Table 3: Results comparison on ImageNet-1k. Our full results with other models are in Appendix M.4.
@@
Transformer DeiT-S [74] 224 12 22M 79.8
DeiT-S + AttnScale [80] 224 12 22M 80.7
DeiT-S + FeatScale [80] 224 12 22M 80.9
DeiT-S + ContraNorm [28] 224 12 22M 80.4
Swin-S [45] 224 12 50M 82.9
DeiT-S [74] 224 24 43M 80.5
CaiT-S [75] 224 24 47M 82.6
DeiT-S + AttnScale [80] 224 24 44M 81.1
DeiT-S + FeatScale [80] 224 24 44M 81.3
DeiT-S + ContraNorm [28] 224 24 43M 80.7
GFSA DeiT-S + GFSA 224 12 22M 81.1
DeiT-S + GFSA 224 24 43M 81.5
CaiT-S + GFSA 224 24 47M 82.8
Swin-S + GFSA 224 12 50M 83.0
5.4 Experiments on Graph-level Tasks
Setting.
To evaluate the efficacy of GFSA on graph-level tasks, we conduct experiments on a broader range of datasets. We use datasets from Long-Range Graph Benchmark (LRGB) [21] (e.g., Peptide-func and Peptide-struct), Benchmarking GNNs [22] (e.g., ZINC, MNIST, CIFAR10), Open Graph Benchmark (OGB) dataset [32] (e.g., Molhiv and MolTox21), and OGB-LSC dataset (i.e., PCQM4M-LSC) [33]. We choose Graphormer [94], Graph-ViT [30], and GPS [63] as our backbone architectures, following their original experimental protocols for fair comparison. For GPS, we replace its self-attention module with our GFSA while maintaining its best configuration and other hyperparameters. For Graph-ViT, we apply GFSA to the Hadamard self-attention method, which He et al. [30] propose as optimal. For a detailed experimental setting, see Appendix O.1.
Results.
Tables 55, and 6 show consistent performance improvements when GFSA is integrated with backbone architectures. Graph-ViT + GFSA shows improvements on all datasets. On Peptide-func, it achieves a 0.65% increase in AP. Notably, in PCQM4M, incorporating GFSA improves the validation MAE by 7.20%. Due to space constraints, the results with standard deviation are included in Appendix O.2.
Table 4: Results on ZINC
Method #Params MAE ($\downarrow$)
Graphormer 500K 0.1240
+ GFSA 500K 0.1189
Table 5: Results on PCQM4M and PCQM4Mv2
Method #Params PCQM4M PCQM4Mv2
Train ($\downarrow$) Validate ($\downarrow$) Train ($\downarrow$) Validate ($\downarrow$)
Graphormer 48.3M 0.0535 0.1286 0.0250 0.0862
+ GFSA 48.3M 0.0312 0.1193 0.0249 0.0860
Table 6: Experimental evalutation of GFSA plugged into GPS and Graph-ViT. Results marked with $\dagger$ indicate settings where we conducted our own experiments due to unavailable Hadamard self-attention performance in He et al. [30]’s paper.
Method Peptide-func Peptide-struct MNIST CIFAR10 Molhiv MolTOX21 ZINC
AP ($\uparrow$) MAE ($\downarrow$) Accuracy ($\uparrow$) Accuracy ($\uparrow$) ROCAUC ($\uparrow$) ROCAUC ($\uparrow$) MAE ($\downarrow$)
GPS 0.6535$\pm$0.0041 0.2500$\pm$0.0005 0.9805$\pm$0.0013 0.7230$\pm$0.0036 0.070$\pm$0.004
+ GFSA 0.6593$\pm$0.0094 0.2496$\pm$0.0013 0.9814$\pm$0.0014 0.7244$\pm$0.0048 0.069$\pm$0.002
Graph-ViT 0.6919$\pm$0.0085 0.2474$\pm$0.0016 0.9820$\pm$0.0005 0.6967$\pm$0.0040 0.7792$\pm$0.0149 0.7851$\pm$0.0077 0.0849$\pm$0.0047
+ GFSA 0.6964$\pm$0.0025 0.2461$\pm$0.0024 0.9826$\pm$0.0004 0.6987$\pm$0.0028 0.7830$\pm$0.0109 0.7895$\pm$0.0069 0.0845$\pm$0.0032